OpenStudy (anonymous):

What is the vertical asymptote, horizontal asymptote, and hole at (x-2)(x+1)/(3x-1)(x-2)

OpenStudy (alexwee123):

there is a discontinuity at x=2, y asymp=y=-1,1/3 and x asymps=x=-1

OpenStudy (anonymous):

I think x=2 is a removable singularity. (x-2) cancels out. Asymptotes; Vertical x =1/3 Horizontal y=1/3